# Quantum Polymorphisms and Commutativity Gadgets

### v1

@@ 1. $(\mathbf{noc_{1}})$
    For any quantum function $Q\colon\{u,v\}\to_{q}A$ such that $Q_{u,a}$ and $Q_{v,b}$ commute for all $a,b\in A$, there exists a quantum non-oracular homomorphism $Q^{\prime}\colon\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ extending $Q$.
@@ 2. $(\mathbf{noc_{2}})$
    For every quantum non-oracular homomorphism $Q\colon\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ and all $a,b\in A$, it holds that $Q_{u,a},Q_{v,b}]=0$.
We now give the non-oracular analogue of Theorem 16. Just like in the oracular setting, the characterisation of commutativity gadgets via quantum polymorphisms (stated in the result below) also holds for the weaker algebraic non-oracular commutativity gadgets of [Zeman], which are defined by replacing ($(\mathbf{noc_{1}})$) in Definition 43 with the weaker condition stated below:
@@ 1. $(\mathbf{noc_{1}^{\prime}})$
    For every $a,b\in A$, there exists a quantum non-oracular homomorphism $Q\colon\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ such that $Q_{u,a}=\operatorname{id}$ and $Q_{v,b}=\operatorname{id}$.
@@
Let $\mathbb{A}$ be a finite structure. The following are equivalent:
  1. (1)
@@
    The proof follows from exactly the same argument as for Proposition 20.
@@
    Take an algebraic non-oracular commutativity gadget $\mathbb{G}$ with distinguished elements $u,v$, a non-oracular quantum polymorphism $Q\colon\mathbb{A}^{n}\operatorname{\to_{qno}}\mathbb{A}$, two tuples $\mathbf{a},\mathbf{b}\in A^{n}$, and two elements $c,d\in A$. We proceed as in Proposition 21. We first build a non-oracular quantum polymorphism $W^{(i)}:\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ for each $i\in[n]$ with the property that $W^{(i)}_{u,a_{i}}=W^{(i)}_{v,b_{i}}=\operatorname{id}$, using ($(\mathbf{noc_{1}^{\prime}})$). Then, we take the tensor product $W=\bigotimes_{i\in[n]}W^{(i)}$ of such homomorphisms, and we define $U:\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ as the composition of $Q$ and $W$. On the one hand, ($(\mathbf{noq_{2}})$) guarantees that $U_{u,c},U_{v,d}]=0$. On the other hand, expanding the composition yields $U_{u,c}=Q_{\mathbf{a},c}\otimes\operatorname{id}$ and $U_{v,d}=Q_{\mathbf{b},d}\otimes\operatorname{id}$. Combining the two facts, we conclude that $Q_{\mathbf{a},c},Q_{\mathbf{b},d}]=0$, and we deduce that $Q$ is classical.∎
The next result shows that non-oracular q-definitions, combined with non-oracular commutativity gadgets, give logspace reductions from $\operatorname{CSP^{*}}$s to $\operatorname{CSP^{**}}$s.
@@
  • for every $R\in\rho$, there exists a non-oracular q-definition for $\mathbb{A}$ and $R^{\mathbb{B}}$.
Then $\operatorname{CSP^{*}}(\mathbb{B})$ reduces to $\operatorname{CSP^{**}}(\mathbb{A})$ in logspace.
@@
Let $\mathbb{X}$ be an instance of $\operatorname{CSP^{*}}(\mathbb{B})$. Recall that $\operatorname{Gaif}(\mathbb{X})$ denotes the Gaifman digraph of $\mathbb{X}$. We define an instance $\mathbb{Y}$ of $\operatorname{CSP^{**}}(\mathbb{A})$ as follows. Let the domain of $\mathbb{Y}$ originally consist of the disjoint union of $X$, together with a copy $\mathbb{G}_{R}^{(\mathbf{x})}$ of $\mathbb{G}_{R}$ for each symbol $R\in\rho$ and each tuple $\mathbf{x}\in R^{\mathbb{X}}$, as well as a copy $\mathbb{H}^{(x,y)}$ of $\mathbb{H}$ for each edge $(x,y)$ of $\operatorname{Gaif}(\mathbb{X})$. For $R\in\rho$ of some arity $r\in\mathbb{N}$, $\mathbf{x}\in R^{\mathbb{X}}$, and $i\in[r]$, we denote by $g_{i}^{(\mathbf{x})}$ the copy of $g_{i}$ in $\mathbb{G}_{R}^{(\mathbf{x})}$. Similarly, we denote by $u^{(x,y)}$ (resp. $v^{(x,y)}$) the copy of $u$ (resp. $v$) in $\mathbb{H}^{(x,y)}$. For every edge $(x,y)$ of $\operatorname{Gaif}(\mathbb{X})$, glue $u^{(x,y)}$ with $x$ and glue $v^{(x,y)}$ with $y$. For each $R\in\rho$ of some arity $r$, each $\mathbf{x}\in R^{\mathbb{X}}$, and each $i\in[r]$, glue $g_{i}^{(\mathbf{x})}$ with $x_{i}$. Call this instance $\mathbb{Y}$.
(Completeness.) Suppose that $Q\colon\mathbb{X}\to_{q}\mathbb{B}$ is a quantum homomorphism. For every symbol $R\in\rho$ of some arity $r\in\mathbb{N}$ and every $\mathbf{x}\in R^{\mathbb{X}}$, consider the $\{R\}$-structure $\mathbb{A}^{\prime}=\mathbb{A}_{R^{\mathbb{B}}}$ with domain $A$ and relation $R^{\mathbb{A}^{\prime}}=R^{\mathbb{B}}$. Recall also the $\{R\}$-structure $\mathbb{R}$ described in Definition 8. Setting $W_{i,a}=Q_{x_{i},a}$ for each $i\in[r]$ and each $a\in A=B$ yields a quantum homomorphism $W\colon\mathbb{R}\to\mathbb{A}^{\prime}$. (Observe that this would not be guaranteed if $Q$ was a non-oracular quantum homomorphism.) Then, ($(\mathbf{noq_{1}})$) guarantees the existence of a non-oracular quantum homomorphism $Q^{(\mathbf{x},R)}\colon\mathbb{G}_{R}^{(\mathbf{x})}\operatorname{\to_{qno}}\mathbb{A}$ such that
@@
$Q^{(x,y)}\colon\mathbb{H}^{(x,y)}\operatorname{\to_{qno}}\mathbb{A}$
extending $Q$ in $x$ and $y$. Overall, this is a non-oracular quantum homomorphism $\mathbb{Y}\operatorname{\to_{qno}}\mathbb{A}$.
(Soundness.) Suppose that $Q\colon\mathbb{Y}\operatorname{\to_{qno}}\mathbb{A}$ is a non-oracular quantum homomorphism. Consider the quantum function $Q^{\prime}:X\to B=A$ obtained by restriction of $Q$ to $X\subseteq Y$. We claim that $Q^{\prime}$ is an (oracular) quantum homomorphism from $\mathbb{X}$ to $\mathbb{B}$. To that end, take a symbol $R\in\rho$ of some arity $r\in\mathbb{N}$ and a tuple $\mathbf{x}\in R^{\mathbb{X}}$. Take $i,j\in[r]$. We claim that the PVMs $Q^{\prime}_{x_{i}}$ and $Q^{\prime}_{x_{j}}$ commute. This is easily true if $x_{i}=x_{j}$ (as in this case the PVMs coincide), so suppose that $x_{i}\neq x_{j}$. In this case, $(x_{i},x_{j})$ is an edge in $\operatorname{Gaif}(\mathbb{X})$. Note now that, by construction of $\mathbb{Y}$, the vertices $x_{i}$, $u^{(x_{i},x_{j})}$, and $g_{i}^{(\mathbf{x})}$ are identified in $Y$. Similarly, the vertices $x_{j}$, $v^{(x_{i},x_{j})}$, and $g_{j}^{(\mathbf{x})}$ are identified in $Y$. Since $Q$ induces a non-oracular quantum homomorphism over $\mathbb{H}^{(x_{i},x_{j})}$, we deduce from ($(\mathbf{noc_{2}})$) that the PVMs $Q^{\prime}_{x_{i}}$ and $Q^{\prime}_{x_{j}}$ commute also in this case. Hence, $Q^{\prime}$ satisfies $($$\mathbf{QH_{2}}$)). We now apply ($(\mathbf{noq_{2}})$) to deduce that the quantum function $U$ defined by $(i,a)\mapsto Q_{g_{i}^{(\mathbf{x})},a}=Q^{\prime}_{x_{i},a}$ is a non-oracular quantum homomorphism $\mathbb{R}\operatorname{\to_{qno}}\mathbb{A}^{\prime}$ (where, as before, $\mathbb{A}^{\prime}=\mathbb{A}_{R^{\mathbb{B}}}$). Consider now a tuple $\mathbf{b}\in A^{r}\setminus R^{\mathbb{B}}$. Since $(1,\dots,r)\in R^{\mathbb{R}}$ and $\mathbf{b}\not\in R^{\mathbb{B}}=R^{\mathbb{A}^{\prime}}$, we conclude that
(12) $\prod_{i\in[r]}Q_{x_{i},b_{i}}=0.$
@@
We can finally complete the proof of the main result of this section.
@@
Since $\operatorname{\oplus Pol}(\mathbb{A})=\operatorname{qPol^{no}}(\mathbb{A})$Theorem 44 guarantees that $\mathbb{A}$ has a non-oracular commutativity gadget. Moreover, we know from $\operatorname{qPol^{no}}(\mathbb{A})\subseteq\operatorname{qPol^{no}}(\mathbb{B})$ and Corollary 42 that all the relations of $\mathbb{B}$ have a q-no-definition in $\mathbb{A}$. Note also that $\mathbb{A}$ and $\mathbb{B}$ have the same domain since $\operatorname{qPol^{no}}(\mathbb{A})\subseteq\operatorname{qPol^{no}}(\mathbb{B})$. Hence, the result directly follows from Proposition 45. ∎
@@
Let $\mathbb{T}$ be a tree structure, and let $x,y$ be elements of $\mathbb{T}$. Let $Q\colon\mathbb{T}\operatorname{\to_{qno}}\mathbb{A}$ be a non-oracular quantum homomorphism. If $Q_{x,a}Q_{y,b}\neq 0$, there exists a homomorphism $h\colon\mathbb{T}\to\mathbb{A}$ such that $h(x)=a$ and $h(y)=b$.
@@
We prove this by induction on the size of $\mathbb{T}$. Assume first that $\mathbb{T}$ consists of a single constraint $(z_{1},\dots,z_{r})\in R^{\mathbb{T}}$ with $x=z_{i}$ and $y=z_{j}$ for some $i,j\in\{1,\dots,r\}$. For the argument below, we can assume without loss of generality that $i<j$ up to swapping $x$ and $y$. Then we have
@@
and since $Q_{x,a}Q_{y,b}\neq 0$, it must be that there exists $(c_{1},\dots,c_{r})\in R^{\mathbb{A}}$ such that $c_{i}=a$ and $c_{j}=b$. We obtain a homomorphism $h\colon\mathbb{T}\to\mathbb{A}$ by setting $h(z_{k})=c_{k}$ for all $k\in\{1,\dots,r\}$.
Otherwise, $\mathbb{T}$ consists of the union of some trees $\mathbb{T}_{1},\dots,\mathbb{T}_{m}$ that share a single vertex $z$. Since $\sum_{c}Q_{x,a}Q_{z,c}Q_{z,c}Q_{y,b}=Q_{x,a}(\sum_{c}Q_{z,c})Q_{y,b}=Q_{x,a}Q_{y,b}\neq 0$, there exists a $c$ such that $Q_{x,a}Q_{z,c}\neq 0$ and $Q_{z,c}Q_{y,b}\neq 0$. By induction hypothesis on the subtrees $\mathbb{T}_{1},\dots,\mathbb{T}_{m}$, there exist homomorphisms $h_{i}\colon\mathbb{T}_{i}\to\mathbb{A}$ such that $h_{i}(z)=c$ for all $i$, and such that if $x\in T_{i}$ (resp. $y\in\mathbb{T}_{i}$) then $h_{i}(x)=a$ (resp. $h_{i}(y)=b$). The union of these homomorphisms gives a homomorphism $h\colon\mathbb{T}\to\mathbb{A}$ with $h(x)=a$ and $h(y)=b$. ∎
@@
Suppose that $Q\colon\mathbb{R}\to_{q}(A;R)$ is a quantum homomorphism. Then we have $Q=h_{1}\oplus\dots\oplus h_{d}$, where each $h_{i}$ is a map $\{1,\dots,r\}\to A$ such that $(h_{i}(1),h_{i}(2))\in R$. Thus, each $h_{i}$ can be extended to a homomorphism $g_{i}\colon\mathbb{G}\to\mathbb{A}$ since $\mathbb{G}$ is a classical gadget for $R$. It follows that $Q^{\prime}:=g_{1}\oplus\dots\oplus g_{d}$ is a (non-oracular) quantum homomorphism $\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ satisfying the conclusion of ($(\mathbf{noq_{1}})$).
To prove ($(\mathbf{noq_{2}})$), let $Q\colon\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ be a non-oracular quantum homomorphism. Let $x_{1},x_{2}$ be the distinguished vertices of $\mathbb{G}$. Suppose that $Q_{x_{1},a_{1}}Q_{x_{2},a_{2}}\neq 0$. Since $\mathbb{G}$ is a tree, there exists by Lemma 46 a homomorphism $h\colon\mathbb{G}\to\mathbb{A}$ such that $h(x_{i})=a_{i}$ for $i\in\{1,2\}$. Thus, $(a_{1},a_{2})\in R$, and therefore the quantum function $Q^{\prime}\colon\{1,2\}\to A$ defined by $Q^{\prime}_{i,a}=Q_{x_{i},a}$ satisfies $($$\mathbf{QH_{1}}$)) and is a non-oracular quantum homomorphism $\mathbb{R}\operatorname{\to_{qno}}(A;R)$. ∎
We conclude this section by showing that also the non-oracular version of the CSP parameterised by odd cycles is undecidable. The result is proved in same way as its oracular counterpart.
Corollary 48.
For every odd $m\geq 3$, $\operatorname{CSP^{**}}(\mathbb{C}_{m})$ is undecidable.

### v2

@@ 1. $(\mathbf{noc_{1}})$
    For any quantum function $Q\colon\{u,v\}\to_{\operatorname{q}}A$ such that $Q_{u,a}$ and $Q_{v,b}$ commute for all $a,b\in A$, there exists a quantum non-oracular homomorphism $Q^{\prime}\colon\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ extending $Q$.
@@ 2. $(\mathbf{noc_{2}})$
    For every quantum non-oracular homomorphism $Q\colon\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ and all $a,b\in A$, it holds that $Q_{u,a},Q_{v,b}]=0$.
Note that, unlike in the oracular case Definition 47, a non-oracular commutativity gadget is not the same as a q-no-definition for $A^{2}$ in $\mathbb{A}$. We now give the non-oracular analogue of Theorem 51. Just like in the oracular setting, the characterisation of commutativity gadgets via quantum polymorphisms (stated in the result below) also holds for the weaker algebraic non-oracular commutativity gadgets of [34], which are defined by replacing ($(\mathbf{noc_{1}})$) in Definition 96 with the weaker condition stated below:
@@ 1. $(\mathbf{noc_{1}^{\prime}})$
    For every $a,b\in A$, there exists a quantum non-oracular homomorphism $Q\colon\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ such that $Q_{u,a}=\operatorname{id}$ and $Q_{v,b}=\operatorname{id}$.
@@
Let $\mathbb{A}$ be a relational structure. The following are equivalent:
  1. 1.
@@
    The proof follows exactly the same argument as for Proposition 53.
@@
    Take an algebraic non-oracular commutativity gadget $\mathbb{G}$ with distinguished elements $u,v$, a non-oracular quantum polymorphism $Q\colon\mathbb{A}^{n}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$, two tuples $\mathbf{a},\mathbf{b}\in A^{n}$, and two elements $c,d\in A$. We proceed as in Proposition 54. We first build a non-oracular quantum polymorphism $W^{(i)}\colon\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ for each $i\in[n]$ with the property that $W^{(i)}_{u,a_{i}}=W^{(i)}_{v,b_{i}}=\operatorname{id}$, using ($(\mathbf{noc_{1}^{\prime}})$). Then, we take the tensor product $W=\bigotimes_{i\in[n]}W^{(i)}$ of such homomorphisms, and we define $U\colon\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ as the composition of $Q$ and $W$. On the one hand, ($(\mathbf{noq_{2}})$) guarantees that $U_{u,c},U_{v,d}]=0$. On the other hand, expanding the composition yields $U_{u,c}=Q_{\mathbf{a},c}\otimes\operatorname{id}$ and $U_{v,d}=Q_{\mathbf{b},d}\otimes\operatorname{id}$. Combining the two facts, we conclude that $Q_{\mathbf{a},c},Q_{\mathbf{b},d}]=0$, and we deduce that $Q$ is classical.∎
The next result shows that non-oracular q-definitions, combined with non-oracular commutativity gadgets, give logspace reductions from $\operatorname{CSP_{q}}$ to $\operatorname{CSP_{q}^{no}}$.
@@
  • $\mathbb{A}$ q-no-defines $\mathbb{B}$.
Then $\operatorname{CSP_{q}}(\mathbb{B})$ reduces to $\operatorname{CSP_{q}^{no}}(\mathbb{A})$ in logspace.
@@
Let $\mathbb{X}$ be an instance of $\operatorname{CSP_{q}}(\mathbb{B})$. Henceforth in this proof, we shall view $\operatorname{Gaif}(\mathbb{X})$ as the directed graph obtained by viewing each edge $\{x,y\}$ as two directed edges $(x,y)$ and $(y,x)$. We define an instance $\mathbb{Y}$ of $\operatorname{CSP_{q}^{no}}(\mathbb{A})$ as follows. Let the domain of $\mathbb{Y}$ originally consist of the disjoint union of $X$, together with a copy $\mathbb{G}_{R}^{(\mathbf{x})}$ of $\mathbb{G}_{R}$ for each symbol $R\in\rho$ and each tuple $\mathbf{x}\in R^{\mathbb{X}}$, as well as a copy $\mathbb{H}^{(x,y)}$ of $\mathbb{H}$ for each edge $(x,y)$ of $\operatorname{Gaif}(\mathbb{X})$. For $R\in\rho$ of some arity $r\in\mathbb{N}$, $\mathbf{x}\in R^{\mathbb{X}}$, and $i\in[r]$, we denote by $g_{i}^{(\mathbf{x})}$ the copy of $g_{i}$ in $\mathbb{G}_{R}^{(\mathbf{x})}$. Similarly, we denote by $u^{(x,y)}$ (resp. $v^{(x,y)}$) the copy of $u$ (resp. $v$) in $\mathbb{H}^{(x,y)}$. For every edge $(x,y)$ of $\operatorname{Gaif}(\mathbb{X})$, glue $u^{(x,y)}$ with $x$ and glue $v^{(x,y)}$ with $y$. For each $R\in\rho$ of some arity $r$, each $\mathbf{x}\in R^{\mathbb{X}}$, and each $i\in[r]$, glue $g_{i}^{(\mathbf{x})}$ with $x_{i}$. Call this instance $\mathbb{Y}$.
(Completeness.) Suppose that $Q\colon\mathbb{X}\to_{\operatorname{q}}\mathbb{B}$ is a quantum homomorphism. For every symbol $R\in\rho$ of some arity $r\in\mathbb{N}$ and every $\mathbf{x}\in R^{\mathbb{X}}$, consider the $\{R\}$-structure $\mathbb{A}^{\prime}=(A;R^{\mathbb{B}})$ with domain $A$ and relation $R^{\mathbb{A}^{\prime}}=R^{\mathbb{B}}$. Recall also the $\{R\}$-structure $\mathbb{R}$ described in Definition 93. Setting $W_{i,a}=Q_{x_{i},a}$ for each $i\in[r]$ and each $a\in A=B$ yields a quantum homomorphism $W\colon\mathbb{R}\to_{\operatorname{q}}\mathbb{A}^{\prime}$. (Observe that this would not be guaranteed if $Q$ was a non-oracular quantum homomorphism.) Then, ($(\mathbf{noq_{1}})$) guarantees the existence of a non-oracular quantum homomorphism $Q^{(\mathbf{x},R)}\colon\mathbb{G}_{R}^{(\mathbf{x})}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ such that
@@
$Q^{(x,y)}\colon\mathbb{H}^{(x,y)}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$
extending $Q$ in $x$ and $y$. Overall, this is a non-oracular quantum homomorphism $\mathbb{Y}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$.
(Soundness.) Suppose that $Q\colon\mathbb{Y}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ is a non-oracular quantum homomorphism. Consider the quantum function $Q^{\prime}:X\to B=A$ obtained by restriction of $Q$ to $X\subseteq Y$. We claim that $Q^{\prime}$ is an (oracular) quantum homomorphism from $\mathbb{X}$ to $\mathbb{B}$. To that end, take a symbol $R\in\rho$ of some arity $r\in\mathbb{N}$ and a tuple $\mathbf{x}\in R^{\mathbb{X}}$. Take $i,j\in[r]$. We claim that the PVMs $Q^{\prime}_{x_{i}}$ and $Q^{\prime}_{x_{j}}$ commute. This is easily true if $x_{i}=x_{j}$ (as in this case the PVMs coincide), so suppose that $x_{i}\neq x_{j}$. In this case, $(x_{i},x_{j})$ is an edge in $\operatorname{Gaif}(\mathbb{X})$. Note now that, by construction of $\mathbb{Y}$, the vertices $x_{i}$, $u^{(x_{i},x_{j})}$, and $g_{i}^{(\mathbf{x})}$ are identified in $Y$. Similarly, the vertices $x_{j}$, $v^{(x_{i},x_{j})}$, and $g_{j}^{(\mathbf{x})}$ are identified in $Y$. Since $Q$ induces a non-oracular quantum homomorphism over $\mathbb{H}^{(x_{i},x_{j})}$, we deduce from ($(\mathbf{noc_{2}})$) that the PVMs $Q^{\prime}_{x_{i}}$ and $Q^{\prime}_{x_{j}}$ commute also in this case. Hence, $Q^{\prime}$ satisfies $($$\mathbf{QH_{2}}$)). We now apply ($(\mathbf{noq_{2}})$) to deduce that the quantum function $U$ defined by $(i,a)\mapsto Q_{g_{i}^{(\mathbf{x})},a}=Q^{\prime}_{x_{i},a}$ is a non-oracular quantum homomorphism $\mathbb{R}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}^{\prime}$, where, as before, $\mathbb{A}^{\prime}=(A;R^{\mathbb{B}})$. Consider now a tuple $\mathbf{b}\in A^{r}\setminus R^{\mathbb{B}}$. Since $(1,\dots,r)\in R^{\mathbb{R}}$ and $\mathbf{b}\not\in R^{\mathbb{B}}=R^{\mathbb{A}^{\prime}}$, we conclude that
$\prod_{i\in[r]}Q_{x_{i},b_{i}}=0.$ (18)
@@
We can finally complete the proof of Theorem 92.
@@
Since $\operatorname{\oplus Pol}(\mathbb{A})=\operatorname{qPol^{no}}(\mathbb{A})$Theorem 97 guarantees that $\mathbb{A}$ has a non-oracular commutativity gadget. Moreover, we know from $\operatorname{qPol^{no}}(\mathbb{A})\subseteq\operatorname{qPol^{no}}(\mathbb{B})$ and Theorem 13 that $\mathbb{A}$ q-no-defines $\mathbb{B}$. Note also that $\mathbb{A}$ and $\mathbb{B}$ have the same domain since $\operatorname{qPol^{no}}(\mathbb{A})\subseteq\operatorname{qPol^{no}}(\mathbb{B})$. Hence, the result directly follows from Proposition 98. ∎
@@
Let $\mathbb{T}$ be a tree structure, and let $x,y$ be elements of $\mathbb{T}$. Let $Q\colon\mathbb{T}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ be a non-oracular quantum homomorphism. If $Q_{x,a}Q_{y,b}\neq 0$, there exists a homomorphism $h\colon\mathbb{T}\to\mathbb{A}$ such that $h(x)=a$ and $h(y)=b$.
@@
We prove this by induction on the size of $\mathbb{T}$. Assume first that $\mathbb{T}$ consists of a single constraint $(z_{1},\dots,z_{r})\in R^{\mathbb{T}}$ with $x=z_{i}$ and $y=z_{j}$ for some $i,j\in[r]$. For the argument below, we can assume without loss of generality that $i<j$ up to swapping $x$ and $y$. Then we have
@@
and since $Q_{x,a}Q_{y,b}\neq 0$, it must be that there exists $(c_{1},\dots,c_{r})\in R^{\mathbb{A}}$ such that $c_{i}=a$ and $c_{j}=b$. We obtain a homomorphism $h\colon\mathbb{T}\to\mathbb{A}$ by setting $h(z_{k})=c_{k}$ for all $k\in[r]$.
Otherwise, $\mathbb{T}$ consists of the union of some trees $\mathbb{T}_{1},\dots,\mathbb{T}_{m}$ that share a single vertex $z$. Since $\sum_{c}Q_{x,a}Q_{z,c}Q_{z,c}Q_{y,b}=Q_{x,a}(\sum_{c}Q_{z,c})Q_{y,b}=Q_{x,a}Q_{y,b}\neq 0$, there exists a $c$ such that $Q_{x,a}Q_{z,c}\neq 0$ and $Q_{z,c}Q_{y,b}\neq 0$. By the induction hypothesis on the subtrees $\mathbb{T}_{1},\dots,\mathbb{T}_{m}$, there exist homomorphisms $h_{i}\colon\mathbb{T}_{i}\to\mathbb{A}$ such that $h_{i}(z)=c$ for all $i$, and such that if $x\in T_{i}$ (resp. $y\in T_{i}$) then $h_{i}(x)=a$ (resp. $h_{i}(y)=b$). The union of these homomorphisms gives a homomorphism $h\colon\mathbb{T}\to\mathbb{A}$ with $h(x)=a$ and $h(y)=b$. ∎
@@
Suppose that $Q\colon\mathbb{R}\to_{\operatorname{q}}(A;R)$ is a quantum homomorphism. Then we have $Q=h_{1}\oplus\dots\oplus h_{d}$, where each $h_{i}$ is a map $r]\to A$ such that $(h_{i}(1),h_{i}(2))\in R$. Thus, each $h_{i}$ can be extended to a homomorphism $g_{i}\colon\mathbb{G}\to\mathbb{A}$ since $\mathbb{G}$ is a classical gadget for $R$. It follows that $Q^{\prime}:=g_{1}\oplus\dots\oplus g_{d}$ is a (non-oracular) quantum homomorphism $\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ satisfying ($(\mathbf{noq_{1}})$).
To prove ($(\mathbf{noq_{2}})$), let $Q\colon\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ be a non-oracular quantum homomorphism. Let $x_{1},x_{2}$ be the distinguished vertices of $\mathbb{G}$. Suppose that $Q_{x_{1},a_{1}}Q_{x_{2},a_{2}}\neq 0$. Since $\mathbb{G}$ is a tree, there exists by Lemma 99 a homomorphism $h\colon\mathbb{G}\to\mathbb{A}$ such that $h(x_{i})=a_{i}$ for $i\in\{1,2\}$. Thus, $(a_{1},a_{2})\in R$, and therefore the quantum function $Q^{\prime}\colon\{1,2\}\to_{\operatorname{q}}A$ defined by $Q^{\prime}_{i,a}=Q_{x_{i},a}$ satisfies $($$\mathbf{QH_{1}}$)) and is a non-oracular quantum homomorphism $\mathbb{R}\operatorname{\to_{\operatorname{qno}}}(A;R)$. ∎
The next result is the non-oracular version of Proposition 43.
Proposition 101.
Let $\mathbb{A},\mathbb{B}$ be structures, and let $R_{A}\subseteq A^{r}$ and $R_{B}\subseteq B^{r}$ be relations that admit a common non-oracular q-definition $\mathbb{G}$ over $\mathbb{A}$ and $\mathbb{B}$, respectively. Then every non-oracular quantum homomorphism $Q\colon\mathbb{A}\operatorname{\to_{\operatorname{qno}}}\mathbb{B}$ is a non-oracular quantum homomorphism $(A;R_{A})\operatorname{\to_{\operatorname{qno}}}(B;R_{B})$.